OpenStudy (anonymous):

Find the equation of the tangent plane of z^2=x^2+y^2 at (1,1,1)

OpenStudy (anonymous):

@phi @amistre64 @inkyvoyd

OpenStudy (phi):

isn't the gradient of the function f(x,y,z)= x^2+y^2-z^2 normal to the curve ? \[ \nabla f \]= N \] evaluated at 1,1,1 the equation of a plane is \[ N \cdot p = b\] where N is the normal and p is <x,y,z> and b is a scalar

OpenStudy (phi):

Here is an example http://www.ltcconline.net/greenl/courses/107/PartDeriv/tanplane.htm

OpenStudy (anonymous):

I'm not sure? Is the formula: f_x(x,y,z)(x-x0)+f_y(x,y,z)(y-y0)+f_z(x,y,z)(z-z0)=0?

OpenStudy (anonymous):

I think you are thinking of the tangent vector as opposed to the tangent plane

OpenStudy (phi):

It's the same thing. Once I get a normal to the plane, I can plug in a point that I know is on the plane to find b gradient of f is (fx, fy, fz) and if the known point is p= (x0,y0,z0) N dot p = fx*x0 + fy*y0 + fz*z0 = b so the equation would be (fx, fy, fz) dot (x,y,z)= fx*x0 + fy*y0 + fz*z0 which can be written as (x-x0)*fx + (y- y0)*fy + (z-z0)*fz =0 or just \[ N \cdot p= b\] where b is as above

OpenStudy (anonymous):

Thanks!